OpenStudy (anonymous):

Hello everybody! Any idea for find the integral of (e^x)(ln x)/(x^2+1) dx

OpenStudy (perl):

is this the question $$ \Large \int \frac{e^x \cdot ln(x) }{x^2 + 1 } dx $$

OpenStudy (anonymous):

Yes! \[\int\limits \frac{e^x \ln x }{ x^2+1 }dx\]

OpenStudy (anonymous):

good luck with that program i really really doubt you are going to find a closed form for that integral

OpenStudy (anonymous):

lets try the wolf

OpenStudy (anonymous):

I really have 4 exercises looking like this, but this is the easiest.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i= \int+\frac{e^x++ln%28x%29+}{x^2+%2B+1+}+dx

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

is that the exact question?

OpenStudy (anonymous):

I tried with Geogebra too but nothing!

OpenStudy (anonymous):

Yes, it's the question.

OpenStudy (anonymous):

i would say no way

OpenStudy (anonymous):

Wait, there's something more

OpenStudy (anonymous):

But derivative is about x and integral is about t, what can be done here?

OpenStudy (perl):

you dont need to know the antiderivative

OpenStudy (perl):

$$ \Large \frac{d}{dx}\int_{a}^{x} f(t)~ dt = f(x) $$

OpenStudy (anonymous):

That's because with t=a we have the derivative of a constant value and with t=x we got the same expression... And now I can see it! Thanks a lot, I must be so tired to work!

OpenStudy (acxbox22):

i wish i could do calculus :(

OpenStudy (acxbox22):

@perl teach me plez...

OpenStudy (perl):

exactly :)

OpenStudy (perl):

acx ok :)

OpenStudy (acxbox22):

another year perl ;) and i will be there waiting for u to teach ;)

OpenStudy (perl):

sure

OpenStudy (perl):

$$ \Large{ \frac{d}{dx}\int_{a}^{x} f(t)~ dt\\ = \frac{d}{dx} [F(x) - F(a)] = \frac{d}{dx}F(x) -\frac{d}{dx}F(a) \\ = F ' (x) - 0 = f(x) } $$